# Motivic interpretations for iterated integrals on some specific algebraic curves

### v2

@@ 5.2. Galois action
Now, we consider the action of the Galois group $\mathrm{Gal}(F/\mathbb{Q})$ on $\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F$ to investigate the structure of $\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F$ in more detail.
The Galois group $\mathrm{Gal}(F/\mathbb{Q})$ acts on the category of mixed Tate motives $\mathrm{MTM}_{R}$ on $R\subseteq F^{\times}\otimes_{\mathbb{Z}}\mathbb{Q}$ induced by the automorphism as a ring of $R$ induced by the Galois action. In addition, the Galois action is induced on $\mathcal{H}_{F,R}$ since $\mathcal{H}_{F,R}$ is the coordinate ring of Tannakian fundamental group of $\mathrm{MTM}_{R}$. Also, we can extend the $\mathrm{Gal}(F/\mathbb{Q})$-action to $\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F$ diagonally. By Proposition 4.9, each element of $\mathcal{H}_{F,R}$ can be written by motivic iterated integrals if $(F,R)=(F_{g},R_{g})$ or $(F,R)=(F_{h},R_{h})$, then we define the Galois action explicitly for motivic iterated integrals.
@@
If $(F,R)=(F_{f},R_{f})$ for $f\in\{g,h\}$, the Galois group $\mathrm{Gal}(F/\mathbb{Q})$ acts on $\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F$. In addition, $\sigma\in\mathrm{Gal}(F/\mathbb{Q})$ acts on the element written by motivic iterated integrals as
$\displaystyle\sigma(\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;a_{1},\dots,a_{k};1)\otimes c)=\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;{}^{\sigma}{a_{1}},\dots,{}^{\sigma}{a_{k}};1)\otimes{}^{\sigma}{c}.$ (144)
Let us look back at Definition 5.4. Recall that we put $G_{f}:=\mathrm{Gal}(F_{f}/\mathbb{Q})$. Then, we see that each element $\mathrm{I}^{\mathfrak{m}}_{f}(\eta_{1}\cdots\eta_{k})\in\mathcal{H}_{F_{f},R_{f}}\otimes_{\mathbb{Q}}F_{f}$ is $G_{f}$-invariant since so is each $\theta_{p}$ $(p=1,\dots,k)$. Then, it is sufficient to investigate the $G_{f}$-invariant part $\left(\mathcal{H}_{F_{f},R_{f}}\otimes_{\mathbb{Q}}F_{f}\right)^{G_{f}}$ of $\mathcal{H}_{F_{f},R_{f}}\otimes_{\mathbb{Q}}F_{f}$ to discuss the motivic structure of iterated integrals on $X_{f}$.
@@ 5.3. The structure of $\left(\mathcal{H}_{F_{f},R_{f}}\otimes_{\mathbb{Q}}F_{f}\right)^{G_{f}}$
The Galois action on $\mathcal{U}_{F,R}$ is commutative with the non-commutative product of $\mathcal{U}_{F,R}$. That fact is shown in the case of $(F,R)=(\mathbb{Q}(\xi_{4}),\langle 2\rangle_{\mathbb{Q}})$ in [14, p.23, Lemma 5.2.1] and we can show similarly if $(F,R)=(\mathbb{Q}(\xi_{6}),\langle 2,3\rangle_{\mathbb{Q}})$.
@@
    $\dim_{\mathbb{Q}}\left(\mathcal{U}_{F_{g},R_{g}}^{(k)}\otimes_{\mathbb{Q}}F_{g}\right)^{G_{g}}=d_{k,g},$
@@
    $\displaystyle\left(\mathcal{U}_{F_{h},R_{h}}^{(k)}\otimes_{\mathbb{Q}}F_{h}\right)^{G_{h}}$ $\displaystyle=\bigoplus_{\begin{subarray}{c}r\geq 0\\ k-r\text{: even}\end{subarray}}\mathrm{Span}_{\mathbb{Q}}\left\{b_{j_{1}}^{(s_{1})}\cdots b_{j_{r}}^{(s_{r})}\cdot\tau^{l}\otimes 1\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \begin{array}[]{c}j_{1},\dots,j_{r}\in\mathbb{Z}_{\geq{1}},l\in\mathbb{Z}_{\geq{0}},\\ s_{i}\in\begin{cases}\{1,2\},&j_{i}=1\\ \{1\},&j_{i}>1\end{cases},\leavevmode\nobreak\ \sum_{i=1}^{r}j_{i}+l=k\end{array}\right\}$ (178)
@@
The following table summarizes the value of $D_{f}^{(k)}$ and conjectural value of $\dim_{\mathbb{Q}}Z_{f}^{(k)}$ for each $f\in\{g,h\}$ and small weight $k\in\mathbb{Z}_{\geq{1}}$ obtained by numerical experiments. The value of each iterated integral was calculated by SageMath, and linear independence was determined by the PARI/GP library of SageMath. As an observation of this table, we can check Corollary 1.3 numerically for $k=0,1,\dots,5$. The table also shows that there is rich relational $\mathbb{Q}$-linear relations along iterated integrals on $Y_{f}$. The difference between $D_{f}^{(k)}$ and $\dim_{\mathbb{Q}}Z_{f}^{(k)}$ is due to the fact that $\mathrm{per}\left(\left(\mathcal{H}_{F_{f},R_{f}}^{(k)}\otimes_{\mathbb{Q}}F_{f}\right)^{\mathrm{Gal}(F_{f}/\mathbb{Q})}\right)\not\subseteq Z_{f}^{(k)}$. Also, if $f=g$, $\dim_{\mathbb{Q}}Z_{g}^{(k)}$ is expected to be $1\displaystyle 2^{k}-$ 1 and the element $u\in\left(\mathcal{H}_{F_{f},R_{f}}^{(k)}\otimes_{\mathbb{Q}}F_{f}\right)^{\mathrm{Gal}(F_{f}/\mathbb{Q})}$ such that $\mathrm{per}(u)\notin Z_{g}^{(k)}$ would be $u=(\log^{\mathfrak{m}}2)^{k}:=\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;-1;1)^{k}$ according to the numerical experiments.
@@
We do not claim that the direct sum decomposition is true, but since Remark 5.2, the equality (215) is true if so is the period conjecture (i.e. $\mathrm{per}:\mathcal{H}_{F_{f},R_{f}}\rightarrow\mathbb{C}$ is injective).
@@ 5.4. Alternating multiple mixed values
Therefore, AMMVs can be written as $\mathbb{Q}(\sqrt{{-1}})$-linear sum of multiple $L$-values of level 44 44. In addition, each differential forms in the equation 224 is $\mathrm{Gal}(\mathbb{Q}(\sqrt{-1})/\mathbb{Q})$-invariant as same as $\theta_{j}$ in the item I. Then, we have the following proposition as an analogue of Theorem 1.2.
@@
For any $\mathbf{k}=(k_{1},\dots,k_{d})\in\left(\mathbb{Z}_{\geq{1}}\right)^{d}$ and $\boldsymbol{\varepsilon}=(\varepsilon_{1},\dots,\varepsilon_{d}),\boldsymbol{\sigma}=(\sigma_{1},\dots,\sigma_{d})\in\{\pm 1\}^{d}$ with $(\epsilon_{d},\sigma_{d})\neq(1,1)$, there exists an element $M_{\boldsymbol{\sigma}}^{\mathfrak{m},\boldsymbol{\varepsilon}}(\mathbf{k})\in\left(\mathcal{H}_{\mathbb{Q}(\sqrt{-1}),\langle 2\rangle_{\mathbb{Q}}}^{(k)}\otimes_{\mathbb{Q}}\mathbb{Q}(\sqrt{-1})\right)^{\mathrm{Gal}(\mathbb{Q}(\sqrt{-1})/\mathbb{Q})}$ such that
@@
$\mathrm{per}\left(\mathcal{H}_{\mathbb{Q}(\sqrt{-1}),\langle 2\rangle_{\mathbb{Q}}}^{(k)}\right)=\mathrm{MLV}_{4}^{(k)}$
@@
If the period conjecture is true (i.e. $\mathrm{per}:\mathcal{H}_{\mathbb{Q}(\sqrt{-1}),\langle 2\rangle_{\mathbb{Q}}}\rightarrow\mathbb{C}$ is injective),
@@
  • [9] M. Kaneko K. Ihara and D. Zagier. Derivation and double shuffle relations for multiple zeta values. Compos. Math., 142:307–338, 2006.

### v3

@@ 5.2. Galois action
In the following, the extension $F/\mathbb{Q}$ is assumed to be Galois. We consider the action of the Galois group $\mathrm{Gal}(F/\mathbb{Q})$ on $\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F$ to investigate the Galois fixed part of $\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F$ in detail.
Let $S$ be a finite set of finite places in $F$ and $\mathcal{O}_{F,S}$ the ring of $S$-integers of $F$. Let $\mathcal{G}_{F}$ be the motivic Galois group associated to the canonical fiber functor of $\mathrm{MT}(F)$ defined in [4, Section 1.2]. It is a pro-algebraic affine group scheme over $\mathbb{Q}$. Let $\mathcal{O}(\mathcal{G}_{F})$ be the ring of the regular functions on $\mathcal{G}_{F}$. The Galois action of $\mathrm{Gal}(F/\mathbb{Q})$ on $F$ induces the action on each object of $\mathrm{MT}(F)$. Further, this action induces the action on $\mathcal{O}(\mathcal{G}_{F})$ and on its subalgebra $\mathcal{H}_{F}$. This action is explicitly computed in [8, p.212, Theorem 1.1] for the motivic iterated integrals belonging to $\mathcal{H}_{F}$. Let $\mathcal{G}^{\mathrm{dR}}_{F}$ be the motivic Galois group associated to the de Rham fiber functor of $\mathrm{MT}(F)$ defined in [4, Section 2.9]. It is a pro-algebraic affine group scheme over $F$. Let $\mathcal{O}(\mathcal{G}^{\mathrm{dR}}_{F})$ be the ring of the regular functions on $\mathcal{G}^{\mathrm{dR}}_{F}$. Then, by [4, Proposition 2.10], $\mathcal{G}^{\mathrm{dR}}_{F}\simeq\mathcal{G}_{F}\times_{\mathrm{Spec}\hskip 1.42262pt\mathbb{Q}}\mathrm{Spec}\hskip 1.42262ptF$ and it yields $\mathcal{O}(\mathcal{G}^{\mathrm{dR}}_{F})\simeq\mathcal{O}(\mathcal{G}_{F})\otimes_{\mathbb{Q}}F$ such that the Galois action of $\mathrm{Gal}(F/\mathbb{Q})$ on $\mathcal{O}(\mathcal{G}^{\mathrm{dR}}_{F})$ translates into the diagonal action on $\mathcal{O}(\mathcal{G}_{F})\otimes_{\mathbb{Q}}F$. Now, by definition, $\mathcal{H}_{F,R}$ is a $\mathrm{Gal}(F/\mathbb{Q})$-stable $\mathbb{Q}$-subalgebra of $\mathcal{O}(\mathcal{G}_{F})$ because the canonical fiber functor on $MT(F)$ is stable under the action of $\mathrm{Gal}(F/\mathbb{Q})$. The Galois group $\mathrm{Gal}(F/\mathbb{Q})$ acts diagonally on the base extension $\mathcal{H}_{F,R}\otimes_{\mathbb{Q}}F$ and this action comes from the Galois action of $\mathrm{Gal}(F/\mathbb{Q})$ on $\mathcal{O}(\mathcal{G}^{\mathrm{dR}}_{F})$ via the above isomorphism. By Proposition 4.9, each element of $\mathcal{H}_{F,R}$ can be written by the motivic iterated integrals if $(F,R)=(F_{g},R_{g})$ or $(F,R)=(F_{h},R_{h})$, then we can compute the Galois action explicitly for the motivic iterated integrals with coefficients in $F$ as follows.
@@
If $(F,R)=(F_{f},R_{f})$ for $f\in\{g,h\}$, each element $\sigma$ of $\mathrm{Gal}(F/\mathbb{Q})$ acts on the motivic iterated integrals with coefficients in $F$ as
$\displaystyle\sigma(\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;a_{1},\dots,a_{k};1)\otimes c)=\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;{}^{\sigma}{a_{1}},\dots,{}^{\sigma}{a_{k}};1)\otimes{}^{\sigma}{c},\ c\in F.$ (146)
Proof ..
It follows from the definition of the action and [8, p.212, Theorem 1.1]. ∎
Let us look back at Definition 5.4. Recall that we put $G_{f}:=\mathrm{Gal}(F_{f}/\mathbb{Q})$.
@@
Proof ..
By definition $\mathrm{I}^{\mathfrak{m}}_{f}(\eta_{1}\cdots\eta_{k}):=\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;\theta_{1}\cdots\theta_{k};1)\in\mathcal{H}_{F_{f},R_{f}}\otimes_{\mathbb{Q}}F_{f}$ and since each $\theta_{p}$ is $G_{f}$-invariant, $\mathrm{I}^{\mathfrak{m}}_{f}(\eta_{1}\cdots\eta_{k})$ is also $G_{f}$-invariant and we have
$\mathrm{I}^{\mathfrak{m}}_{f}(\eta_{1}\cdots\eta_{k})\in\left(\mathcal{H}_{F_{f},R_{f}}^{(k)}\otimes_{\mathbb{Q}}F_{f}\right)^{G_{f}}.$
Then, it is sufficient to investigate the $G_{f}$-invariant part $\left(\mathcal{H}_{F_{f},R_{f}}\otimes_{\mathbb{Q}}F_{f}\right)^{G_{f}}$ of $\mathcal{H}_{F_{f},R_{f}}\otimes_{\mathbb{Q}}F_{f}$ to discuss the motivic structure of iterated integrals on $X_{f}$.
@@ 5.3. The structure of $\left(\mathcal{H}_{F_{f},R_{f}}\otimes_{\mathbb{Q}}F_{f}\right)^{G_{f}}$
The Galois action on $\mathcal{U}_{F,R}$ is commutative with the non-commutative product of $\mathcal{U}_{F,R}$. That fact is shown in the case of $(F,R)=(\mathbb{Q}(\xi_{4}),\langle 2\rangle_{F_{g}})$ in [14, p.23, Lemma 5.2.1] and we can show similarly if $(F,R)=(\mathbb{Q}(\xi_{6}),\langle 2,3\rangle_{F_{h}})$.
@@
    $\dim_{\mathbb{Q}}\left(\mathcal{U}_{F_{g},R_{g}}^{(k)}\otimes_{\mathbb{Q}}F_{g}\right)^{G_{g}}=D^{(k)}_{g},$
@@
    $\displaystyle\left(\mathcal{U}_{F_{h},R_{h}}^{(k)}\otimes_{\mathbb{Q}}F_{h}\right)^{G_{h}}$ (172)
    $\displaystyle=\bigoplus_{\begin{subarray}{c}r\geq 0\\ k-r\text{: even}\end{subarray}}\mathrm{Span}_{\mathbb{Q}}\left\{b_{j_{1}}^{(s_{1})}\cdots b_{j_{r}}^{(s_{r})}\cdot\tau^{l}\otimes 1\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \begin{array}[]{c}j_{1},\dots,j_{r}\in\mathbb{Z}_{\geq{1}},l\in\mathbb{Z}_{\geq{0}},\\ s_{i}\in\begin{cases}\{1,2\},&j_{i}=1\\ \{1\},&j_{i}>1\end{cases},\leavevmode\nobreak\ \sum_{i=1}^{r}j_{i}+l=k\end{array}\right\}$ (175)
@@
The above table summarizes the value of $D_{f}^{(k)}$ and conjectural value of $\dim_{\mathbb{Q}}Z_{f}^{(k)}$ for each $f\in\{g,h\}$ and small weight $k\in\mathbb{Z}_{\geq{1}}$ obtained by numerical experiments. The value of each iterated integral was calculated by SageMath, and linear independence was determined by the PARI/GP library of SageMath. As an observation of this table, we can check Corollary 1.3 numerically for $k=0,1,\dots,5$. The table also shows that there is rich relational $\mathbb{Q}$-linear relations along iterated integrals on $Y_{f}$. The difference between $D_{f}^{(k)}$ and $\dim_{\mathbb{Q}}Z_{f}^{(k)}$ is due to the fact that $\mathrm{per}\left(\left(\mathcal{H}_{F_{f},R_{f}}^{(k)}\otimes_{\mathbb{Q}}F_{f}\right)^{\mathrm{Gal}(F_{f}/\mathbb{Q})}\right)\not\subseteq Z_{f}^{(k)}$. Also, if $f=g$, $\dim_{\mathbb{Q}}Z_{g}^{(k)}$ is expected to be $1\displaystyle 2^{k}-$ 1 and the element $u\in\left(\mathcal{H}_{F_{f},R_{f}}^{(k)}\otimes_{\mathbb{Q}}F_{f}\right)^{\mathrm{Gal}(F_{f}/\mathbb{Q})}$ such that $\mathrm{per}(u)\notin Z_{g}^{(k)}$ would be $u=(\log^{\mathfrak{m}}2)^{k}:=\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;-1;1)^{k}$ according to the numerical experiments.
@@
Proof ..
We only show (I). Since $\mathrm{I}^{\mathfrak{m}}_{g}(\eta_{1}\cdots\eta_{k}):=\mathrm{I}^{\mathfrak{m}}_{\mathrm{dch}}(0;\theta_{1}\cdots\theta_{k};1)$ for
$\displaystyle\begin{split}\theta_{j}:=\begin{cases}e_{0}-e_{\xi_{4}}-e_{\xi_{4}^{-1}},&\eta_{j}=\omega_{0},\\ -2e_{1}+e_{\xi_{4}}+e_{\xi_{4}^{-1}},&\eta_{j}=\omega_{1},\\ \xi_{4}^{-1}(e_{\xi_{4}}-e_{\xi_{4}^{-1}}),&\eta_{j}=\omega_{2},\\ e_{0},&\eta_{j}=\omega_{3},\end{cases}\end{split}$
$\mathrm{I}^{\mathfrak{m}}_{g}(\eta_{1}\cdots\eta_{k})$ can be written as a $\mathbb{Q}$-linear sum of $u_{i}\otimes 1$ for $u_{i}\in\mathcal{H}_{\mathbb{Q}(\sqrt{-1}),\langle 2\rangle_{F_{g}}}^{(k)}$ if #$\#\left\{p\leavevmode\nobreak\ |\leavevmode\nobreak\ \eta_{p}=\omega_{2}\right\}$ is even, and $\mathrm{I}^{\mathfrak{m}}_{g}(\eta_{1}\cdots\eta_{k})$ can be written as a $\mathbb{Q}$-linear sum of $u_{i}\otimes\sqrt{-1}$ for $u_{i}\in\mathcal{H}_{\mathbb{Q}(\sqrt{-1}),\langle 2\rangle_{F_{g}}}^{(k)}$ if #$\#\left\{p\leavevmode\nobreak\ |\leavevmode\nobreak\ \eta_{p}=\omega_{2}\right\}$ is odd. The upper bound $D_{g,\delta}^{(k)}$ of the dimension $\dim_{\mathbb{Q}}Z_{g,\delta}^{(k)}$ is given by counting the number of pairs $(j_{1},\dots,j_{r},l)$ satisfying each condition. (II) can be shown similarly. ∎
@@
We do not claim that the direct sum decomposition is true, but by Remark 5.2, equality (202) is true if the period conjecture holds (i.e. $\mathrm{per}:\mathcal{H}_{F_{f},R_{f}}\rightarrow\mathbb{C}$ is injective).
@@ 5.4. Alternating multiple mixed values
Therefore, AMMVs can be written as $\mathbb{Q}(\sqrt{{-1}})$-linear sum of multiple $L$-values of level 44 44. In addition, each differential form in the equation 211 is $\mathrm{Gal}(\mathbb{Q}(\sqrt{-1})/\mathbb{Q})$-invariant as same as $\theta_{j}$ in the item I. Then, we have the following proposition as an analogue of Theorem 1.2.
@@
For any $\mathbf{k}=(k_{1},\dots,k_{d})\in\left(\mathbb{Z}_{\geq{1}}\right)^{d}$ and $\boldsymbol{\varepsilon}=(\varepsilon_{1},\dots,\varepsilon_{d}),\boldsymbol{\sigma}=(\sigma_{1},\dots,\sigma_{d})\in\{\pm 1\}^{d}$ with $(\epsilon_{d},\sigma_{d})\neq(1,1)$, there exists an element $M_{\boldsymbol{\sigma}}^{\mathfrak{m},\boldsymbol{\varepsilon}}(\mathbf{k})\in\left(\mathcal{H}_{\mathbb{Q}(\sqrt{-1}),\langle 2\rangle_{F_{g}}}^{(k)}\otimes_{\mathbb{Q}}\mathbb{Q}(\sqrt{-1})\right)^{\mathrm{Gal}(\mathbb{Q}(\sqrt{-1})/\mathbb{Q})}$ such that
@@
Proof ..
As mentioned above, each AMMV can be written as $\mathbb{Q}(\sqrt{{-1}})$-linear sum of multiple $L$-values of level 44 44. Now, each multiple $L$-value of level 44 44 has a motivic interpretation as an element of $\mathcal{H}_{\mathbb{Q}(\sqrt{-1}),\langle 2\rangle_{F_{g}}}$ Then, there exists an element $M_{\boldsymbol{\sigma}}^{\mathfrak{m},\boldsymbol{\varepsilon}}(\mathbf{k})\in\mathcal{H}_{\mathbb{Q}(\sqrt{-1}),\langle 2\rangle_{F_{g}}}^{(k)}\otimes_{\mathbb{Q}}\mathbb{Q}(\sqrt{-1})$ such that
In addition, since each differential form in the equation 211 is $\mathrm{Gal}(\mathbb{Q}(\sqrt{-1})/\mathbb{Q})$-invariant, then $M_{\boldsymbol{\sigma}}^{\mathfrak{m},\boldsymbol{\varepsilon}}(\mathbf{k})\in\left(\mathcal{H}_{\mathbb{Q}(\sqrt{-1}),\langle 2\rangle_{F_{g}}}^{(k)}\otimes_{\mathbb{Q}}\mathbb{Q}(\sqrt{-1})\right)^{\mathrm{Gal}(\mathbb{Q}(\sqrt{-1})/\mathbb{Q})}$. In particular, since $\dim_{\mathbb{Q}}\left(\mathcal{H}_{\mathbb{Q}(\sqrt{-1}),\langle 2\rangle_{F_{g}}}^{(k)}\otimes_{\mathbb{Q}}\mathbb{Q}(\sqrt{-1})\right)^{\mathrm{Gal}(\mathbb{Q}(\sqrt{-1})/\mathbb{Q})}=2^{k}$ and
$\mathrm{AMMV}^{(k)}\subseteq\mathrm{per}\left(\left(\mathcal{H}_{\mathbb{Q}(\sqrt{-1}),\langle 2\rangle_{F_{g}}}^{(k)}\otimes_{\mathbb{Q}}\mathbb{Q}(\sqrt{-1})\right)^{\mathrm{Gal}(\mathbb{Q}(\sqrt{-1})/\mathbb{Q})}\right),$
we have
$\displaystyle\dim_{\mathbb{Q}}\mathrm{AMMV}^{(k)}\leq 2^{k}.$ (216)
@@
$\mathrm{per}\left(\mathcal{H}_{\mathbb{Q}(\sqrt{-1}),\langle 2\rangle_{F_{g}}}^{(k)}\right)=\mathrm{MLV}_{4}^{(k)}$
@@
If the period conjecture is true (i.e. $\mathrm{per}:\mathcal{H}_{\mathbb{Q}(\sqrt{-1}),\langle 2\rangle_{F_{g}}}\rightarrow\mathbb{C}$ is injective),
@@
  • [9] K. Ihara, M. Kaneko, and D. Zagier. Derivation and double shuffle relations for multiple zeta values. Compos. Math., 142:307–338, 2006.